An ordinary (fair) die is a cube with the numbers 1 through 6 on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment.
Compute the probability of each of the following events.
Event A: The sum is greater than 9.
Event B: The sum is divisible by 2 or 3 (or both).
Round your answer to 2 decimal places.
(a). P(A) =
(b.) P(B) =

Answer by mananth(16946)

Here the total possible outcomes = 36 n(S) =36
Event (A) (4,6),(5,5),(5,6), (6,4),(6,5),(6,6) n(A) = 6
P(A) = n(A)/n(S) = 6/36 = 1/6
Event B: The sum is divisible by 2 or 3 (or both).
There are 24 possible outcomes n(B)=24
P(B)= n(B)/n(S)= 24/36= 2/3
Convert to percent if required
.